solve the following system of equations
x +3y =9 (1)
x=5-3y (2)

Found 2 solutions by ankor@dixie-net.com, checkley77:
Answer by ankor@dixie-net.com(22740)

x + 3y = 9 (1)
x = 5 - 3y (2)
;
Add 3y to both sides of the 2nd equation and you have:
x + 3y = 9
x + 3y = 5
:
Obviously, there is no solution to this system

Answer by checkley77(12844)

x+3y=9 (1)
x=5-3y (2)
x+3y=5
-(x+3y=9) subtract
--------------------
0+0=-4
No unique solution for x or y.